Rς = 20 ΚΩ 5 kΩ - 10 V + υ Λ 8 ΚΩ 2 ΚΩς 5V 1-10V 27 ΚΩξυ, Ενέ ε11ε , και το P5.26_10ed Given: Va+2V Find vo ΤΗ V
Answer #1

given operational amplifier Re = 2062 5K2 JOV V+ t re + 8K2 - LOV 27k2 5V + EARS to ll = 2V Va Aegume given operational ampliV Apply Nodal Analysis at node V-V V_-VO + o 522 2012 V-2 + V-V ZOK 5K (: V = V Virtual shot exist ) V_ - Jx IV1-2 1-V + 5K 206 - 5K + = 20K VO 20K -4+) Vo zok 20k vo 2 20K -3 zok vo = -3V